OpenStudy (aroub):

sqrt[4]{6} times sqrt[4]{36} times sqrt[4]{192}

OpenStudy (aroub):

lol, more of a fail -.- So, this is the question: \[\sqrt[4]{6} \times \sqrt[4]{36} \times \sqrt[4]{192}\]

OpenStudy (aroub):

I know it looks easy, I just keep on getting different answer :/

OpenStudy (aroub):

O.o

OpenStudy (aroub):

I got \[2\sqrt[4]{2592}\] Yeah -.- It looks wrong -.-

hero (hero):

If it has the same root....

hero (hero):

rewrite it like this: \[\large \sqrt[4]{6 \times 36 \times 192}\]

OpenStudy (aroub):

Yes, this is what I did..

OpenStudy (phi):

I would never multiply these numbers together. on the contrary, you want to factor them all are under the root sign. so you have \[ 6 \cdot 6^2 \cdot 6\cdot 2^5 \] or \[ 6^4 \cdot 2^4 \cdot 2\] pull out the powers of 4

OpenStudy (anonymous):